# A Hybrid Quantum Neural Network for Split Learning

### v1

@@ 4.2.2 HQSL Model Variants
We propose two variants of HQSL designed for multivariate binary and multi-class single-channel image dataset classification. Our approach involves constructing an HQNN and subsequently dividing it at the split point (see Fig. 6) so that the initial part of the HQNN represents the client side (classical neural network layers) and the remaining segment functions as the server side (hybrid quantum neural network layers).
@@ 4.2.3 Hybrid Quantum Split Learning Model Training
For the classical layers of HQSL, gradients are computed from the loss function, $\mathcal{L}$, using the classical backpropagation algorithm via the chain rule. However, it becomes non-trivial to do the same for the quantum layer. Schuld et al. devised the parameter-shift rule, which permits backpropagation through the quantum layer [27]. We briefly discuss the main ideas in the following.
Consider a quantum circuit, $f_{q}:\mathbf{Z}\mapsto\mathbf{E}$ ($f_{q}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{N}$), where $f_{q}$ is parameterized by a set of $k$ free parameters, $\mathbf{\Theta}=\{\theta_{1},\theta_{2},\ldots,\theta_{k}\}$, and maps input, $\mathbf{Z}=(z_{1},z_{2},\ldots,z_{M})$ to the measurement outputs, $\mathbf{E}=(e_{1},e_{2},\ldots,e_{N})$. In HQSL, the quantum layer can be considered as being sandwiched between two sets of classical layers: the client-side model, $f_{c}:\mathbf{X}\mapsto\mathbf{Z}$ $(f_{c}:\mathbb{R}^{in}\rightarrow\mathbb{R}^{M})$, and the server-side classical layers, $f_{s}:\mathbf{E}\mapsto\hat{\mathbf{Y}}$ $(f_{s}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{out})$, where $\hat{\mathbf{Y}}$ is the output of HQSL.
@@
In summary, for the quantum layer, during forward propagation, the input classical features are fed to the quantum layer, and the quantum circuit is evolved for a default of 1000 shots. At the end of the quantum circuit evolutions, an expectation value for the measurement is calculated as per Eq. 2. These calculated values are then used as classical features for the next classical layer. During backpropagation, to compute the derivatives with respect to parameter $\theta_{i}\in\{\theta_{1},\theta_{2},\ldots,\theta_{k}\}$ and inputs $z_{i}\in\{z_{1},z_{2},\ldots,z_{M}\}$, two evaluations of the circuit are computed with a $\pm$ shift in the parameter $\theta_{i}$. These two computations are then added according to Eq. 5 and Eq. 6. This is repeated until the derivatives with respect to all $\theta_{i}$ and $z_{j}$ are computed. Using the computed derivatives, we update the gate parameters, $\mathbf{\Theta}$, accordingly. The derivative $\frac{\delta{f_{q}}}{\delta{z_{j}}}$ is used to compute the gradient with respect to the quantum layer input $z_{j}$ as $\frac{\delta{\mathcal{L}}}{\delta{z_{j}}}$ using the chain rule. This gradient is transmitted across the split layer to the client-side model to continue the backpropagation of the loss function, $\mathcal{L}$, to the client-side classical layers. We summarize the HQSL training in Algorithm 1. In the algorithm, $\frac{\delta{(\cdot)}}{\delta{(\cdot)}}$ represents the Jacobian matrix that stores the element-wise gradients of the vector in the numerator with respect to the vector in the denominator.
@@ 4.3 Hybrid Quantum Split Learning with Multiple Clients
Scaling HQSL to accommodate a larger number of clients enables multiple clients to leverage the quantum resources at the central server. The method we use to scale HQSL with K clients follows the basic round-robin protocol and is as follows. We initialize the model parameters of the client side and server side as $\mathbf{W^{C}}$ and $(\mathbf{\Theta},\mathbf{W}^{S})$ respectively. $\mathbf{\Theta}$ corresponds to the quantum layer trainable gate parameters that lie on the server side. We randomly choose a client k ($\textsf{k}\in\{0,1,2,\dotsc,\textsf{K}-1\}$) and train it in collaboration with the server. This consists of one round of forward and backward propagation and makes 1 local epoch for client k. The updated client k model weights $\mathbf{W}^{C^{\prime}}$ are then sent to the next client ($\textsf{k}^{\prime}$) that updates its model weights before another round of forward and backward propagation with the server. This marks the end of client $\textsf{k}^{\prime}$’s local epoch. After all K clients have been served in that 1 global epoch, we move to the next global epoch and resume training client k.
@@ 5.1 Hybrid Quantum Split Learning versus Classical Split Learning: Single Client Experiments
First, we constructed the quantum circuit shown in Fig. 2(b) using PennyLane’s default-qubit device and converted it to a quantum layer. We then constructed the 2 HQSL variants introduced in Section 4.2.2. We built HQSL variant 1 for training on Botnet DGA and Breast Cancer datasets and HQSL variant 2 on MNIST, Fashion-MNIST, and Speech Command spectrogram datasets. In HQSL variant 2, we set the number of output nodes to $n=10$ (10 classes) for the MNIST and Fashion-MNIST datasets and $n=2$ (2 classes) for the Speech Commands spectrograms dataset. We paired each of these 5 experiments with their classical counterparts to benchmark the performance of our HQSL model variants.
For our hyperparameters, we utilized the Adam optimizer with a learning rate of $10^{-3}$ across all experiments, except for the Speech Commands dataset, where a learning rate of $10^{-4}$ provided better convergence. We used the binary cross entropy loss function for classifying the 2 multi-variate datasets and the cross entropy loss function for the 3 image classification tasks. We trained our models as per Algorithm 1 for 100 epochs for Botnet DGA and Breast Cancer datasets and 50 epochs for MNIST, FMNIST, and Speech Commands spectrograms datasets. We report the mean and standard deviation of the accuracies and F1-scores over five folds of the datasets in Fig. 7. Across all pairs of experiments, we set the random seed to be a constant and fixed the training and testing dataset for each of the 5 folds. These tests were also performed on the centralized (unsplit) versions of the split learning models, and the results are reported next.
@@
We present the results of our tests on HQSL and SL for a single client in Fig. 7. We highlight that the test results for the split and centralized HQSL versions are identical.
@@
Although marginal, improvements in both mean accuracies and F1-scores of HQSL were obtained on the Botnet DGA (mean accuracy improvement: $\uparrow 0.1\%$, mean F1-score improvement: $\uparrow 0.4\%$), Breast Cancer ($\uparrow 0.6\%$, $\uparrow 0.7\%$) and MNIST ($\uparrow 0.4\%$, $\uparrow 0.4\%$) datasets. Significant outperformance in terms of mean accuracy and F1-score were obtained on the Fashion-MNIST ($\uparrow 3\%$, $\uparrow 4\%$) and Speech Commands ($\uparrow 1.5\%$, $\uparrow 1.5\%$) datasets. For all of the five datasets used in this study, HQSL proved feasible as it at least matched the testing performance of SL in terms of accuracies and F1-scores. As demonstrated by the results on the Fashion-MNIST and Speech Commands spectrograms datasets, HQSL has the potential to achieve higher accuracy and F1-score compared to its classical counterpart. It is noteworthy to highlight that with just the introduction of 1 quantum layer consisting of a small 2-qubit quantum circuit, HQSL can slightly outperform classical split learning, showcasing the potential power of quantum computation beyond the NISQ era.
@@
It is noteworthy that simulating a quantum computer using a classical computer is computationally intensive, despite having only 1 quantum layer consisting of a simple quantum circuit on the server side of HQSL. The training time for HQSL was significantly longer than that of SL, depending on the size of the dataset. For instance, 5-fold training for 100 epochs, each with 569 data points from the Breast Cancer dataset split in the train:test ratio of 4:1 took approximately 1 hour to run on HQSL. On the other hand, 5-fold cross-validation with 6000 image-label pairs from the MNIST dataset took around 9 hours to complete 50 training epochs of HQSL. In contrast, their classical analogues only took a few minutes to train which is significantly faster than HQSL.
@@
In the single client case, HQSL with a quantum layer consisting of only a small quantum circuit is feasible and offers tangible improvements in classification accuracy and F1-score compared to SL.
@@ 5.2 Hybrid Quantum Split Learning versus Classical Split Learning: Multiple-client Experiments
HQSL accommodates multiple clients without losing its performance significantly. The performance trend is similar to that of the SL counterpart.
6 Strengthening the Security of HQSL against Reconstruction Attacks
In this section, we address the second research question RQ 2 stated in Section 3. Specifically, we investigate the susceptibility of HQSL to reconstruction attacks using smashed data on the server side. First, we introduce the threat model, outlining the potential risks and vulnerabilities of HQSL to reconstruction attacks. Then, we present the reconstruction attack models we consider in this work to recreate a client’s raw input data. These models are designed to exploit the smashed data on the server side.
Considering the rotational properties of the encoding gates in our quantum layer, we propose a noise-based defense mechanism in the form of a Laplacian noise layer. This layer adds randomness to the smashed data before they are transmitted to the server-side model, making it more difficult for the reconstruction attack models to recreate private input raw data. Next, we conduct experiments under different noise parameters to extract the best-performing noise layer whereby HQSL has a clear advantage over SL in two key aspects: (i) impairing the performance of the reconstruction attack models and (ii) maintaining a high classification performance despite the presence of the noise defense layer. In our experiments, we investigate reconstruction attacks considering the image datasets introduced in Section 5, specifically, MNIST, Fashion-MNIST, and the spectrograms from the Speech Commands datasets.

### v2

@@ 4.2.2 HQSL Model Variants
We propose two variants of HQSL designed for multivariate binary and multi-class single-channel image dataset classification. Our approach involves constructing an HQNN and subsequently dividing it at the split point (see Fig. 6) so that the initial part of the HQNN represents the client side (classical neural network layers) and the remaining segment functions as the server side (hybrid quantum neural network layers). This explicit division or splitting of the HQNN into client-side classical layers and server-side hybrid quantum-classical layers constitutes the key novelty of HQSL, distinguishing it from existing HQNN approaches that do not address resource-constrained deployment or collaborative training frameworks.
We assume here that the clients, e.g., IoT devices, have sufficient classical resources to perform client-side computations and lack quantum resources, as stated in Section 3. This is a fair assumption because typical IoT devices have sufficient processing capabilities to execute lightweight classical neural network models, but generally lack the resources and infrastructure for quantum computations. The client-side model portion extracts compressed feature representations (smashed data), to transmit low-dimensional data to the server-side model portion, reducing communication overhead and enabling a shallow quantum circuit in the quantum layer for further processing on the server side.
@@ 4.2.3 Hybrid Quantum Split Learning Model Training
In our HQSL framework, the training process follows a client-server paradigm where data is processed in two distinct stages. The client side, which consists of a classical neural network, processes the raw input data $\mathbf{X}$ and extracts feature representations. These extracted features, referred to as the smashed data $\mathbf{Z}$, are then transmitted to the server side, together with the true labels, $\mathbf{Y}$.
The server side consists of a hybrid quantum-classical sub-model, where, first, the quantum layer processes $\mathbf{Z}$, and outputs measurement expectation values $\mathbf{E}$. $\mathbf{E}$ is now input to the remaining layers of the server-side model portion which, in our HQSL model, are all classical dense layers with $\hat{\mathbf{\mathbf{Y}}}$ as the final output. This structured split ensures that the client operates solely with classical resources, while the server-side quantum component allows the client to benefit from quantum processing without requiring direct access to quantum hardware. The placement of the quantum layer as the first layer of the server-side model portion is a deliberate design choice aimed at enhancing resistance to reconstruction attacks in split learning, as further discussed in Section 6. Next, we outline the computation and backpropagation of gradients through the classical and quantum layers of HQSL to enable effective training.
For the classical layers of HQSL, gradients are computed from the loss function, $\mathcal{L(\mathbf{Y},\hat{\mathbf{Y}})}$, using the classical backpropagation algorithm via the chain rule. However, it becomes non-trivial to do the same for the quantum layer. Schuld et al. devised the parameter-shift rule, which permits backpropagation through the quantum layer [31]. We briefly discuss the main ideas in the following.
Consider a quantum circuit, $f_{q}:\mathbf{Z}\mapsto\mathbf{E}$ ($f_{q}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{N}$), where $f_{q}$ is parameterized by a set of $k$ free parameters, $\mathbf{\Theta}=\{\theta_{1},\theta_{2},\ldots,\theta_{k}\}$, and maps input, $\mathbf{Z}=(z_{1},z_{2},\ldots,z_{M})$ to the measurement outputs, $\mathbf{E}=(e_{1},e_{2},\ldots,e_{N})$. In HQSL, the quantum layer can be considered as being sandwiched between two sets of classical layers: the client-side model layers, $f_{c}:\mathbf{X}\mapsto\mathbf{Z}$ $(f_{c}:\mathbb{R}^{in}\rightarrow\mathbb{R}^{M})$, and the server-side classical layers, $f_{s}:\mathbf{E}\mapsto\hat{\mathbf{Y}}$ $(f_{s}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{out})$, where $\hat{\mathbf{Y}}$ is the output of HQSL. The smashed data $\mathbf{Z}$ is then transferred to the server side to continue forward propagation, together with the true labels $\mathbf{Y}$, to compute the loss $\mathcal{L}$.
@@
In summary, on the server side, for the quantum layer, during forward propagation, the smashed data $\mathbf{Z}$ are fed as classical features to the quantum layer, and the quantum circuit is evolved for a default of 1000 shots. At the end of the quantum circuit evolutions, an expectation value for the measurement is calculated as per Eq. 2. These calculated values are then used as classical features for the next classical layer. During backpropagation, to compute the derivatives with respect to parameter $\theta_{i}\in\{\theta_{1},\theta_{2},\ldots,\theta_{k}\}$ and inputs $z_{i}\in\{z_{1},z_{2},\ldots,z_{M}\}$, two evaluations of the circuit are computed with a $\pm$ shift in the parameter $\theta_{i}$. These two computations are then added according to Eq. 5 and Eq. 6. This is repeated until the derivatives with respect to all $\theta_{i}$ and $z_{j}$ are computed. Using the computed derivatives, we update the gate parameters, $\mathbf{\Theta}$, accordingly. The derivative $\frac{\delta{f_{q}}}{\delta{z_{j}}}$ is used to compute the gradient with respect to the quantum layer input $z_{j}$ as $\frac{\delta{\mathcal{L}}}{\delta{z_{j}}}$ using the chain rule. This gradient is transmitted across the split layer to the client-side model to continue the backpropagation of the loss function, $\mathcal{L}$, to the client-side classical layers. We summarize the HQSL training in Algorithm 1. In the algorithm, $\frac{\delta{(\cdot)}}{\delta{(\cdot)}}$ represents the Jacobian matrix that stores the element-wise gradients of the vector in the numerator with respect to the vector in the denominator.
@@ 4.3 Hybrid Quantum Split Learning with Multiple Clients
Scaling HQSL to accommodate a larger number of clients enables multiple clients to leverage the quantum resources at the central server. The method we use to scale HQSL with K clients follows the basic round-robin protocol and is as follows. We initialize the model parameters of the client side and server side as $\mathbf{W^{C}}$ and $(\mathbf{\Theta},\mathbf{W}^{S})$ respectively. $\mathbf{\Theta}$ corresponds to the quantum layer trainable gate parameters that lie on the server side. We randomly choose a client k ($\textsf{k}\in\{0,1,2,\dotsc,\textsf{K}-1\}$) and train it in collaboration with the server. This consists of one round of forward and backward propagation and makes 1 local epoch for client k. The updated client k model weights $\mathbf{W}^{C^{\prime}}$ are then sent to the next client ($\textsf{k}^{\prime}$) that updates its model weights before another round of forward and backward propagation with the server. This marks the end of client $\textsf{k}^{\prime}$’s local epoch. After all K clients have been served in that 1 global epoch, we move to the next global epoch and resume training client k. It is important to highlight that during each communication round between a client and the server, only low-dimensional data is transmitted, as the output layer of the client-side model consists of only 3 nodes, as shown in Fig. 6. This significantly reduces the communication overhead, especially when considering a very large number of clients, e.g., IoT devices in edge networks. A detailed analysis or optimization of communication efficiency is beyond the scope of this work. We leave a thorough investigation of strategies to further minimize communication overhead in HQSL for future research.
@@ 5.1 Hybrid Quantum Split Learning versus Classical Split Learning: Single Client Experiments
First, we constructed the quantum circuit shown in Fig. 2(b) using PennyLane’s default-qubit device and converted it to a quantum layer. To build HQSL variants 1 and 2 as introduced in Section 4.2.2, we first constructed their respective client-side and server-side model portions separately, as illustrated in Fig. 6, ensuring that each variant adhered to the split learning framework. We emphasize again here that the quantum layer was incorporated as the initial layer of the server-side model portion of each HQSL variant. This separation allows for the collaborative training of the client-side and server-side model portions in HQSL while maintaining the distinct roles of the client and the server as per the split learning protocol.
We built HQSL variant 1 for training on Botnet DGA and Breast Cancer datasets and HQSL variant 2 on MNIST, Fashion-MNIST, and Speech Command spectrogram datasets. In HQSL variant 2, we set the number of output nodes to $n=10$ (10 classes) for the MNIST and Fashion-MNIST datasets and $n=2$ (2 classes) for the Speech Commands spectrograms dataset. We paired each of these 5 experiments with their classical counterparts to benchmark the performance of our HQSL model variants.
For our hyperparameters, we utilized the Adam optimizer with a learning rate of $10^{-3}$ across all experiments, except for the Speech Commands dataset, where a learning rate of $10^{-4}$ provided better convergence. We used the binary cross entropy loss function for classifying the 2 multi-variate datasets and the cross entropy loss function for the 3 image classification tasks. We trained our HQSL model variants as per Algorithm 1 for 100 epochs for Botnet DGA and Breast Cancer datasets and 50 epochs for MNIST, FMNIST, and Speech Commands spectrograms datasets. We report the mean and standard deviation of the accuracies and F1-scores over five folds of the datasets in Fig. 7. Across all pairs of experiments, we set the random seed to be a constant and fixed the training and testing datasets for each of the 5 folds. These experiments were also performed on the centralized (unsplit) versions of each split learning model. We highlight here that the training process of the centralized version of HQSL is identical to that of a single-node HQNN. We next compare the classification performance results of single-client HQSL against their classical SL equivalent, as discussed in the next section.
Hardware and Noisy Simulator Experiments. To evaluate the feasibility of deploying HQSL on real or noisy quantum hardware, we tested the model on both real IBMQ devices and simulated noisy environments using qiskit_aer [8] with increasing single and two-qubit depolarizing noise levels ($p=0.05$ to $0.09$). We reference calibrated error rates from the 127-qubit ibm_brisbane backend, where median single-qubit gate errors (SX) are typically below $3e^{-4}$, and two-qubit gate (ECR) errors have a median of $7e^{-3}$, which are substantially below the noise levels used in our noisy simulation experiments. These simulations allow us to test the limits of HQSL’s robustness under exaggerated noise conditions that exceed those observed on current hardware. We benchmarked against Pennylane’s noise-free default.qubit device. Due to the high execution time associated with accessing real quantum hardware, we limited our evaluation to a single fold of the Botnet DGA dataset. Specifically, we trained single-client HQSL using the training set on the default.qubit simulator and evaluated its performance on the various backends using the corresponding testing set. These results are presented in Fig. 8.
@@
We present the results of our tests on HQSL and SL for a single client in Fig. 7. We highlight that the test results for the split and centralized HQSL versions are identical, confirming that the splitting process does not degrade model accuracy and F1-score. This demonstrates that HQSL successfully maintains the benefits of split learning (collaborative training of a resource-constrained client with a quantum-enhanced server) while achieving the same performance as its centralized single-node HQNN counterpart.
@@
Although marginal, improvements in both mean accuracies and F1-scores of HQSL were obtained on the Botnet DGA (mean accuracy improvement: $\uparrow 0.1\%$, mean F1-score improvement: $\uparrow 0.4\%$), Breast Cancer ($\uparrow 0.6\%$, $\uparrow 0.7\%$) and MNIST ($\uparrow 0.4\%$, $\uparrow 0.4\%$) datasets. Significant outperformance in terms of mean accuracy and F1-score were obtained on the Fashion-MNIST ($\uparrow 3\%$, $\uparrow 4\%$) and Speech Commands ($\uparrow 1.5\%$, $\uparrow 1.5\%$) datasets. The visually narrow box plot for the Speech Commands dataset in the HQSL case is a result of the low variance across the five cross-validation folds. This indicates the higher stability of our HQSL model’s performance on this dataset compared to the other datasets and to SL’s performance of the Speech Commands dataset. For completeness, we provide the performance results for the Speech Commands dataset in Appendix F.
@@
For all of the five datasets used in this study, HQSL proved feasible as it at least matched the testing performance of SL in terms of accuracies and F1-scores. As demonstrated by the results on the Fashion-MNIST and Speech Commands spectrograms datasets, HQSL has the potential to achieve higher accuracy and F1-score compared to its classical counterpart. We highlight that with just the introduction of 1 quantum layer consisting of a small 2-qubit quantum circuit, HQSL can slightly outperform classical split learning, showcasing the potential power of quantum computation beyond the NISQ era.
@@
It is noteworthy that simulating a quantum computer using a classical computer is computationally intensive, despite having only 1 quantum layer consisting of a simple quantum circuit on the server side of HQSL. The training time for HQSL was significantly longer than that of SL, depending on the size of the dataset. For instance, 5-fold training for 100 epochs, each with 569 data points from the Breast Cancer dataset split in the train:test ratio of 4:1 took approximately 1 hour to run on HQSL. On the other hand, 5-fold training with 6000 image-label pairs from the MNIST dataset took around 9 hours to complete 50 training epochs of HQSL. In contrast, their classical analogues only took a few minutes to train which is significantly faster than HQSL.
Evaluation Results on Real Quantum Devices and Noisy Simulators. To assess the practical viability of HQSL under real-world noise, we evaluated the model on both real IBMQ backends and noisy simulators. As shown in Fig. 8, HQSL achieves consistently high accuracy and F1-scores on ibm_sherbrooke, ibm_strasbourg, and ibm_brussels, similar to the noise-free default.qubit simulator, indicating that currently available noisy quantum hardware is sufficient to support HQSL deployment with our proposed qubit-efficient data-loading quantum circuit in its quantum layer. Interestingly, significant performance drops occurred only when using unrealistically high depolarizing noise models in simulation (e.g., $p=0.08$ or $p=0.09$), suggesting that these levels represent overly pessimistic device conditions. These results further support the feasibility of deploying HQSL on real hardware.
Figure 8: Deployment accuracy and F1-score of HQSL evaluated on the Botnet DGA dataset across multiple backends. default.qubit represents HQSL’s noise-free simulation results, which match the performance obtained on IBMQ hardware (ibm_sherbrooke, ibm_strasbourg, ibm_brussels). Noisy simulators (qiskit_aer) with increasing depolarizing noise $p$ demonstrate a drop in performance as noise levels increase. This shows HQSL’s noise-resilience during deployment on currently available hardware
@@
In the single client case, HQSL with a quantum layer consisting of only a small quantum circuit is feasible and offers tangible improvements in classification accuracy and F1-score compared to SL. Moreover, evaluations on real quantum hardware and noisy simulators highlight HQSL’s practical feasibility and robustness for deployment on currently available quantum devices.
@@ 5.2 Hybrid Quantum Split Learning versus Classical Split Learning: Multiple-client Experiments
By leveraging the scalability property of split learning, HQSL can accommodate multiple clients without dropping performance. The performance trend is similar to that of the SL counterpart.
6 Enhancing the Resistance of HQSL against Reconstruction Attacks
In this section, we address the second research question RQ 2 stated as follows: How do we strengthen such hybrid quantum split learning schemes against data privacy leakage and reconstruction attacks? Specifically, we investigate the susceptibility of HQSL to reconstruction attacks using smashed data on the server side. First, we introduce the threat model, outlining the potential risks and vulnerabilities of HQSL to reconstruction attacks. Then, we present the reconstruction attack models we consider in this work to recreate a client’s raw input data. These models are designed to exploit the smashed data on the server side.
Considering the rotational properties of the encoding gates in our quantum layer, we propose a noise-based defense mechanism in the form of a Laplacian noise layer. The split structure of HQSL conveniently allows us to insert a noise layer at the client-server boundary, directly on the transmitted smashed data. Specifically, this layer adds controlled randomness to the smashed data before they are transmitted to the server-side model, making it more difficult for the reconstruction attack models to recreate private input raw data. Next, we conduct experiments under different noise parameter settings to extract the best-performing noise layer whereby HQSL has a clear advantage over SL in two key aspects: (i) impairing the performance of the reconstruction attack models and (ii) maintaining a high classification performance despite the presence of the noise defense layer. In our experiments, we investigate reconstruction attacks considering the image datasets introduced in Section 5, specifically, MNIST, Fashion-MNIST, and the spectrograms from the Speech Commands datasets.